OpenStudy (anonymous):

how do you find points on a parabola (that aren't the zeros) without a graphing calculator?

OpenStudy (anonymous):

@iambatman @Miracrown @Kainui @SolomonZelman @triciaal

OpenStudy (triciaal):

use the value of x in the function and the result is y value

OpenStudy (anonymous):

........? im function is f(x)=-x^5-3x^4+21x^3+47x^2-120x-144

OpenStudy (triciaal):

y = f(x) for a parabola f(x) = ax^2 + bx + c

OpenStudy (anonymous):

do i just plug in random numbers for x?

OpenStudy (triciaal):

example f(2) means x = 2 y = f(x)=-x^5-3x^4+21x^3+47x^2-120x-144 f(2)= -(2)^5-3*(2)^4 +21*(2)^3 + 47*(2)^2-120*(2)- 144=

OpenStudy (anonymous):

oh okay! so i continue doing that until i find a whole number

OpenStudy (triciaal):

when you plug in a value for x and the answer is 0 there is no remainder and that value is a factor of the polynomial

OpenStudy (triciaal):

this is the same function you had earlier and the zeros were -4, -1 and 3 so when you use these values for x the answer is 0

OpenStudy (anonymous):

ok. Thank you!